OpenStudy (anonymous):

Find the smallest positive fraction x/y in lowest terms with x and y as integers so that when 7 is added to both the numerator/denominator of the fraction, the value of this new fraction equals the value of the original fraction (x/y) when 2 is subtracted from the numerator and 8 is subtracted from the denominator.

OpenStudy (anonymous):

or in other words, \[\frac{x+7}{y+7}=\frac{x-2}{y-8}\] find x/y with x and y as integers

OpenStudy (anonymous):

anyone?